how to find the length of a diagonal if you know the area=100 m2
Answer by CharlesG2(834)

how to find the length of a diagonal if you know the area=100 m2
A area of rectangle = L length * W width = 100 m^2
use Pythagorean Theorem to solve for D diagonal
D^2 = L^2 + W^2 (can put L in terms of W or W in terms of L, lets use W in terms of L)
D^2 = L^2 + (100/L)^2 and if in terms of W then:
D^2 = W^2 + (100/W)^2
this would be really easy if this were a square because then the diagonal would be one of the sides multiplied by the square root of 2, in this case 10*sqrt(2)